OpenStudy (anonymous):

Find the area of a parallelogram if a base and corresponding altitude have the indicated lengths. Base 11/2 feet, altitude 6 inches. 3/4 sq. ft. 1 1/4 sq. ft. 4 1/2 sq. ft.

OpenStudy (anonymous):

11/2 * 1/2 = 11/4 = 2 3/4 sq ft

OpenStudy (anonymous):

maybe first one shld be 2 3/4 ?

OpenStudy (phi):

I would guess 11/2 should be 1 and 1/2 or 3/2 ft.

OpenStudy (vishweshshrimali5):

jimmy is right

OpenStudy (anonymous):

yeah it was suppsed to be 1 1/2